# BioProBench: Comprehensive Dataset and Benchmark in Biological Protocol Understanding and Reasoning

### v1

@@ Abstract
Biological protocols are fundamental to reproducible and safe life science research. While LLMs excel on general tasks, their systematic evaluation on these highly specialized, accuracy-critical, and inherently procedural texts remains limited. In this work, we present BioProBench—the first large-scale, integrated multi-task benchmark for biological protocol understanding and reasoning. While limited benchmarks have touched upon specific aspects like protocol QA, BioProBench provides a comprehensive suite of five core tasks: Protocol Question Answering, Step Ordering, Error Correction, Protocol Generation, and Protocol Reasoning, enabling a holistic evaluation of LLMs on procedural biological texts. Built upon 27K original protocols, it yields nearly 556K high-quality structured instances. We evaluate 12 mainstream open/closed-source LLMs on BioProBench. Experimental results reveal that while top models preform well on surface understanding tasks (e.g., $\sim$70PQA-Acc., $>$64% ERR F1), struggle significantly with deep reasoning and structured generation tasks like ordering and generation. Furthermore, model comparisons reveal diverse performance: certain open-source models approach closed-source levels on some tasks, yet bio-specific small models lag behind general LLMs, indicating limitations on complex procedural content. Overall, our findings underscore that procedural reasoning within biological protocols represents a significant challenge for current LLMs. BioProBench serves as a standardized framework to diagnose these specific limitations and guide the development of AI systems better equipped for safely automating complex scientific procedures. The code and data are available at: https://github.com/YuyangSunshine/bioprotocolbench and https://huggingface.co/datasets/GreatCaptainNemo/BioProBench. footnotetext: Equal Contribution: $\star$; Corresponding author: $\dagger$.
@@ 1 Introduction
Large language models (LLMs) have demonstrated significant progress across general biomedical text mining tasks, ranging from document classification and named-entity recognition to relation extraction, summarization, and question answering lievin2024can ; nori2023can . While specialized models like BioBERT lee2020biobert , BioGPT luo2022biogpt , and BiomedGPT luo2023biomedgpt show domain-specific adaptation through fine-tuning on biomedical corpora, they typically lack the inherent architecture or training specifically geared towards handling complex procedural information like that found in protocols. Consequently, existing benchmarks for biomedical text processing, including BioASQ tsatsaronis2015overview , PubMedQA jin2019pubmedqa , LAB-Bench laurent2024lab and BixBench mitchener2025bixbench , despite driving progress in areas like question answering and data interpretation, have not systematically addressed the unique challenges of interpreting, reasoning about, and generating biological protocols—tasks central to experimental science.
@@
To fill this gap, we propose Biological Protocol Benchmark (BioProBench), the first large-scale, integrated multi-task benchmark for biological protocol understanding and reasoning, specifically designed for LLMs (Figure 1). BioProBench features a layered data design to support diverse model development stages: a raw corpus of 27K protocols for pretraining or RAG; a substantial downstream training set of over 550K structured instances across five fine-grained tasks for adaptation; and a held-out test set of 1,000 examples per task for standardized benchmarking. This design allows models to first leverage broad procedural knowledge from the raw corpus, then specialize on task nuances with the large training set, and finally be rigorously evaluated on consistent, representative examples. To accurately quantify procedural knowledge, we complement standard NLP metrics (e.g., BLEU, METEOR) with novel domain-specific measures: keyword-based content metrics and embedding-based structural metrics. Evaluating 12 state-of-the-art LLMs (including open-source, closed-source, and domain-specific models) on BioProBench reveals fundamental limitations in their current ability to interpret and reason about biological protocols. While advanced models like Gemini-2.5-pro-exp google2025gemini2_5_pro_exp achieved 70.27% ACC. on PQA, and 65% F1 on ERR, performance drops significantly on tasks requiring deeper procedural understanding and structured generation, with ordering (ORD-EM $\sim$ 50%) and generation (GEN-BLEU $<$ 15%) posing major challenges. Model comparisons reveal diverse performance: certain open-source models (e.g., Deepseek-R1 guo2025deepseek and Deepseek-V3 liu2024deepseekv3 ) approach closed-source levels on some tasks, yet smaller, domain-specific models often underperform due to insufficient capacity for capturing complex procedural dependencies. These findings provide crucial insights and directions for developing next-generation biomedical AI systems capable of professional-level experimental reasoning.
@@
  • We introduce BioProBench, the first large-scale, integrated multi-task benchmark for biological protocol understanding and reasoning, featuring over 556K high-quality structured instances derived from 27K real-world protocols.
  • We design a comprehensive set of five tasks, grounded in common ambiguities, safety hazards, and nested dependencies intrinsic to real-world biological experiments, challenging LLMs on diverse procedural understanding and generation capabilities.
@@ 2.1 Data Collection and Processing
As demonstrated in Figure. 1 (a),we collected 26,933 full-text protocols from 6 authoritative online resources: Bio-protocol, Protocol Exchange, JOVE, Nature Protocols, Morimoto Lab, and Protocols.io. The URLs for these resources are provided in Appendix D. These protocols span 16 biological subfields. While Cell Biology and Imaging are predominant, the collection includes examples from Genomics, Immunology, Synthetic Biology, and more, ensuring broad domain coverage, as shown in Figures 1 (b) and (c). This broad distribution reflects the interdisciplinary nature of modern biological research and enhances the benchmark’s generalizability across diverse experimental contexts.
@@
Figure 2: Overview of the BioProBench dataset construction workflow. The process begins with collecting and structuring raw protocols. This is followed by LLM-based data augmentation and an LLM-in-the-loop process to generate diverse task instances. Finally, tasks undergo rigorous quality filtering through self-filtering mechanisms and expert review from domain specialists to ensure accuracy and relevance.
@@ 2.2 Task Instance Generation
Based on the enhanced data, we defined five core tasks: protocol question answering (QA), step ordering (ORD), error correction (ERR), protocol generation (GEN), and protocol reasoning (REA). These tasks are designed to provide fine-grained data for both training/fine-tuning and evaluation, addressing language ambiguities and operational risks inherent in real biological experiments, thereby facilitating LLMs’ development and robust benchmarking in this domain. Overall, BioProBench comprises 556,171 structured task instances. Their distribution and generation methods are summarized below and illustrated in Figure 1 (d). Specifically, 380,697 instances originate from publicly licensed resources (Bio-protocol, Protocol Exchange, protocols.io) and will be released under the CC BY 4.0 license, while approximately 175,474 instances are derived from restricted-access protocols (JOVE, Morimoto Lab, Nature Protocols) and are used for internal evaluation only.
@@
Protocol Question Answering (PQA) task simulates common information retrieval scenarios in biological experiments by querying three critical dimensions: reagent dosages, parameter values, and operational instructions. PQA addresses real-world ambiguities in protocol texts, including inconsistent units, undefined concentration ranges, and unordered procedural steps. Using Deepseek-V2 liu2024deepseek , we automatically construct multiple-choice questions with 5 options for each scenario segment and carefully designed the perturbation options to realistically reproduce the distractors that may be encountered in the laboratory. Three questions were generated for each dimension, introducing reasonable distractors and covering high-risk ambiguities. The final dataset contains 79,869 questions, with distractors calibrated to reflect both syntactic and semantic ambiguities prevalent in laboratory workflows. Details on the PQA format are in Appendix C.1.
Error Correction (ERR) task assesses the ability to identify and correct errors related to safety and result risks caused by text ambiguity or potential input mistakes. ERR focuses on typical deviations in reagents, parameters, and operations, simulating high-risk scenarios like volume overrides, incorrect concentrations, temperature conflicts, or reagent mislabeling. Using Deepseek-V3 liu2024deepseekv3 , we subtly modified key locations in original protocol steps to create error examples, generating an equal number of correct counterparts. The ERR task comprises 157,882 instances (wrong-correct pairings),and it can evaluate and improve the ability to locate subtle errors in large amounts of text and apply biological knowledge to make safe and accurate corrections. Details on format are in Appendix C.3.
@@
Step Ordering(ORD) task aims to enhance the deep understanding of protocol hierarchy and procedural dependencies. Given the nested nature of biological protocols, we design two ordering forms: Top-Step (ordering main stages) and Child-Step (ordering sub-steps within a stage). The ORD task includes 79,098 instances (comprising 23,171 Top-Step and 55,927 Child-Step instances based on internal breakdown). Instances are derived from original protocol steps and shuffled according to predefined rules (without using LLMs for shuffling) to create challenging ordering problems. Through the dual challenges of the global and local levels, the ORD task can accurately evaluate the model’s ability to restore the experimental temporal logic, identify causal relationships, and maintain the consistency of the protocol structure. The detail format is described in Appendix C.2.
@@
Protocol Generation (GEN) task aims to enhance and evaluate the model’s ability to follow instructions under long text creation and professional constraints. To address challenges like incomplete descriptions and inconsistent terminology in real protocols, we extract key information from enhanced data and use Deepseek-R1 guo2025deepseek to automatically generate three tasks of different difficulty levels: Easy (atomic steps, no dependencies), Standard (two-level nesting, basic dependencies) and Difficult (three or more levels, complex dependencies). The examples generated by LLM must be strictly extracted from the original text rather than constructed. Up to five easy, three standard, and one difficult task are generated per protocol, retaining original step numbering. Priority is given to representative subtasks if source material is insufficient. The GEN task contains 183,392 instances. Details on the format are in Appendix C.4.
@@
Protocol Reasoning (REA) task introduces Chain of Thought (CoT) prompting based on ERR and GEN tasks to probe explicit reasoning pathways regarding experimental intent and potential risks. For GEN task, we designed a structured CoT template comprising <Objective>, <Precondition>, <Phase>, <Prameter>and <Structure>. This data guides the model to sort out the experimental objectives and conditional dependencies, ensuring that the generated content is both scientifically reasonable and in line with the field format specifications. For ERR task, CoT instances describe identified error types and analyze potential experimental risks and safety hazards. Based on the above two sets of CoT templates, the REA task constructs about 78,824 reasoning instances. Details on the CoT format are in Appendix C.5.
@@ 2.3 Multi-stage quality control
To guarantee data reliability and quality, we implemented a multi-stage quality control process. This process begins with a three-phase automated self-filtering pipeline. First, automated checks validated structural integrity (e.g. correct step indices, answer–question alignment). Second, template-based and semantic-similarity tests removed any mismatches between generated and reference answers. Third, outputs for challenging cases were regenerated and the best versions selected. Following automated filtering, a randomly selected 10% of instances underwent manual review by an expert group comprising over 5 PhD-level biologists. This manual review focused on critical aspects such as experimental safety constraints, accuracy of technical terminology, and the scientific validity of procedural steps and reasoning chains. This multi-stage approach effectively balances dataset scale with the high fidelity required for evaluating complex biological protocol comprehension and generation.
@@ 3 Evaluation Metrics
For each task, we evaluate model performance using a combination of standard and domain-specific metrics. Table 1 provides an overview of all metrics used across different tasks. Details on standard metrics and LLM-as-a-Judge consistency are provided in the Appendix E. Below, we detail the novel domain-specific metrics we introduce: keyword-based content metrics and embedding-based structural metrics.
@@ 4 Experiments
We conducted experiments with a diverse range of LLMs to provide a broad perspective on current capabilities for biological protocol tasks. To this end, we selected 12 state-of-the-art models, encompassing three main categories: Closed-Source LLMs (o3-mini openai2025o3mini , GPT-4o achiam2023gpt , GPT-4-turbo openai2023gpt4turbo , Claude-3.7-sonnet anthropic2025claude37sonnet , Gemini-2.0-flash google2024gemini2 , Gemini-2.5-pro-exp google2025gemini2_5_pro_exp ), Open-Source LLMs (Deepseek-R1 guo2025deepseek , DeepSeek-V3 liu2024deepseek , QwQ-32B qwen2025qwq32b , Qwen-2.5-72b-instruct qwen2024qwen2_5_72b ), and Domain-Specific Bio-LLMs (BioMedGPT-10B luo2023biomedgpt , BioMistral-7B labrak2024biomistral , BioMistral-7B-DARE labrak2024biomistral ). For evaluation, we constructed a held-out test set consisting of approximately 1,000 representative samples per task. Full implementation details are provided in Appendix F.
@@
4.1 Performance on Protocol Question Answering (PQA)
On PQA task, evaluated LLMs show varied proficiency in comprehending and extracting information from biological protocols (Table 2). The top performance is achieved by the closed-source model Gemini-2.5-pro-exp (70.27%Acc.). However, leading open-source models, notably Deepseek-R1 (67.83%Acc.) and Deepseek-V3 (66.58%Acc.), demonstrate strong capabilities, surpassing several other closed-source models. This highlights the increasing viability of advanced open-source models for domain-specific comprehension tasks, significantly narrowing the performance gap with closed-source alternatives in this domain. In contrast, the dedicated domain-specific models, BioMedGPT-10B and BioMistral-7B, performed exceptionally poorly (20-21%Acc.), exhibiting very high failure rates (33-53%Failed). This suggests that their limited parameter count and primary optimization for general biomedical text are insufficient for the broader procedural and technical understanding required by biological protocols, a capability currently better captured by larger, generalist LLMs. Accuracy generally correlated positively with confidence calibration, as measured by the Brier Score (BS). The top performer, Gemini-2.5-pro-exp, showed the best calibration (0.1949 BS), closely followed by Deepseek-R1 (0.1965 BS). Conversely, models with lower accuracy, such as GPT-4-turbo (57.92% Acc.), exhibited poorer calibration (0.3403 BS). This consistency suggests that models with stronger comprehension capabilities also provide more reliable confidence estimates for their predictions.

### v2

@@ Abstract
Biological protocols are fundamental to reproducibility and safety in life science research. While large language models (LLMs) perform well on general tasks, their systematic evaluation on these highly specialized, accuracy-critical, and inherently procedural texts remains limited. In this work, we present BioProBench, the first large-scale, multi-task benchmark for biological protocol understanding and reasoning. While there are several benchmark tasks involving protocol question answering, BioProBench provides a comprehensive suite of five core tasks: Protocol Question Answering, Step Ordering, Error Correction, Protocol Generation, and Protocol Reasoning, enabling a holistic evaluation of LLMs on procedural biological texts. Built upon 27K original protocols, it yields nearly 556K high-quality structured instances. We evaluate 12 mainstream open/closed-source LLMs. Experimental results reveal that some models perform well on basic understanding tasks (e.g., $\sim$70PQA-Acc., $>$64% ERR F1), but struggle significantly with deep reasoning and structured generation tasks like ordering and generation. Furthermore, model comparisons show diverse performance: certain open-source models approach closed-source levels on some tasks, yet bio-specific small models lag behind general LLMs, indicating limitations on complex procedural content. Overall, BioProBench, through its task design and experimental findings, systematically reveals the fundamental challenges for current LLMs in procedural knowledge understanding, deep adaptability to specific domains, reliability of structured reasoning, and handling of sophisticated precision and safety constraints, providing key directions for future AI in the field of scientific experiment automation. The code and data are available at: https://github.com/YuyangSunshine/bioprotocolbench and https://huggingface.co/datasets/BioProBench/BioProBench. footnotetext: Equal Contribution: $\star$; Corresponding author: $\dagger$.
@@ 1 Introduction
Large language models (LLMs) have demonstrated significant progress across general biomedical text mining tasks nori2023can ; thirunavukarasu2023large ; lievin2024can ; yang2024advancing ; singhal2025toward , ranging from document classification and named-entity recognition to relation extraction, summarization, and question answering. While specialized models like BioBERT lee2020biobert , BioGPT luo2022biogpt , and BioMedGPT luo2023biomedgpt show domain-specific adaptation through fine-tuning on biomedical corpora, they typically lack the inherent architecture or training specifically geared towards handling complex procedural information like that found in protocols. Consequently, existing benchmarks for biomedical text processing, including BioASQ tsatsaronis2015overview , PubMedQA jin2019pubmedqa , LAB-Bench laurent2024lab and BixBench mitchener2025bixbench , despite driving progress in areas like question answering and data interpretation, have not systematically addressed the unique challenges of interpreting, reasoning about, and generating biological protocols, tasks central to experimental science.
@@
To fill this gap, we propose Biological Protocol Benchmark (BioProBench), the first large-scale, integrated multi-task dataset and benchmark for biological protocol understanding and reasoning, specifically designed for LLMs (Figure 1). BioProBench features a layered data design to support diverse model development stages: a raw corpus of 27K protocols for pretraining or RAG; a substantial downstream training set of over 550K structured instances across five fine-grained tasks for adaptation; and a held-out test set of 1,000 examples per task for benchmarking. This design provides both high-quality original corpora and focus on domain-specific fine-grained tasks. To accurately quantify procedural knowledge, we complement standard NLP metrics with novel domain-specific measures: keyword-based content metrics and embedding-based structural metrics. Evaluating 12 state-of-the-art LLMs (including open-source, closed-source, and domain-specific models) on BioProBench reveals fundamental limitations in their current ability to interpret, reason about and generate biological protocols. While advanced models like Gemini-2.5-pro-exp google2025gemini2_5_pro_exp achieves 70.27% PQA-ACC., and 65% ERR-F1, performance drops significantly on tasks requiring deeper procedural understanding (ORD-EM $\sim$ 50%) and structured generation(GEN-BLEU $<$ 15%). Furthermore, in ERR and GEN tasks, zero-shot and one-shot chain of thought (CoT) prompts are set to guide the correct reasoning process and improve performance. In addition, some open-source models approach closed-source levels on certain tasks, while smaller, domain-specific models often underperform due to insufficient capacity for capturing complex procedural dependencies. The fine-grained task settings and evaluation results provide important insights and directions for developing biomedical foundation models capable of professional-level experimental reasoning.
@@
  • We introduce BioProBench, the first large-scale, integrated multi-task dataset and benchmark for biological protocol understanding and reasoning, featuring over 556K high-quality structured instances derived from 27K high-quality protocols.
  • We design a comprehensive set of five tasks that challenge LLMs in different aspects of procedural comprehension and generation capabilities based on ambiguities, safety hazards, and nested dependencies commonly found in real-world biological experiments.
@@ 2.1 Data Collection and Processing
As demonstrated in Figure. 1 (a),we collected 26,933 full-text protocols from 6 authoritative resources: Bio-protocol, Protocol Exchange, JOVE, Nature Protocols, Morimoto Lab, and Protocols.io. The details for these resources are provided in Appendix D. These protocols span 16 biological subfields. While Cell Biology and Imaging are predominant, the collection includes examples from Genomics, Immunology, Synthetic Biology, and more, ensuring broad domain coverage, as shown in Figures 1 (b) and (c). This broad distribution reflects the interdisciplinary nature of modern biological research and enhances the benchmark’s generalizability across diverse experimental contexts.
@@
Figure 2: Workflow of the BioProBench dataset construction. The process begins with collecting and structuring raw protocols and LLM-based data augmentation. This is followed by LLM-in-the-loop process to generate diverse task instances. Finally, tasks quality filtering through Self-Filtering mechanisms and Expert Feedback from domain specialists to ensure accuracy and relevance.
@@ 2.2 Task Instance Generation
Based on the enhanced data, we defined five core tasks: protocol question answering (QA), step ordering (ORD), error correction (ERR), protocol generation (GEN), and protocol reasoning (REA). These tasks are designed to provide fine-grained data for both training/fine-tuning and evaluation, addressing language ambiguities and operational risks inherent in real biological experiments, thereby facilitating LLMs’ development and robust benchmarking in this domain. Overall, BioProBench comprises 556,171 structured task instances. Their distribution and samples are illustrated in Figure 1 (d) and 3. Specifically, 380,697 instances originate from publicly licensed resources (Bio-protocol, Protocol Exchange, protocols.io) and will be released, while approximately 175,474 instances are derived (details descripted in Appendix D).
Figure 3: Samples across each tasks from BioProBench.
@@
Protocol Question Answering (PQA) task simulates common information retrieval scenarios in biological experiments by querying three critical dimensions: reagent dosages, parameter values, and operational instructions. PQA addresses real-world ambiguities in protocol texts, including inconsistent units, undefined concentration ranges, and unordered procedural steps. Using Deepseek-V2 liu2024deepseek , we automatically construct multiple-choice questions with 5 options for each scenario segment and carefully designed the perturbation options to realistically reproduce the distractors that may be encountered in the laboratory. Three questions were generated for each dimension, introducing reasonable distractors and covering high-risk ambiguities. Detail format is in AppendiC.1.
@@
Step Ordering(ORD) task aims to enhance the deep understanding of protocol hierarchy and procedural dependencies. Given the nested nature of biological protocols, we design two ordering forms: Top-Step (ordering main stages) and Child-Step (ordering sub-steps within a stage) to construct the dual challenges of the global and local levels. Instances are derived from original protocol steps and shuffled according to predefined rules (without using LLMs for shuffling) to create challenging ordering problems. The detail format is described in Appendix C.2.
Error Correction (ERR) task assesses the ability to identify and correct errors related to safety and result risks caused by text ambiguity or potential input mistakes. ERR focuses on typical deviations in reagents, parameters, and operations, simulating high-risk scenarios like volume overrides, incorrect concentrations, temperature conflicts, or reagent mislabeling. Using Deepseek-V3 liu2024deepseekv3 , we subtly modified key locations in original protocol steps to create error examples, generating an equal number of correct counterparts. Details on the format are in Appendix C.3.
@@
Protocol Generation (GEN) task aims to enhance and evaluate the model’s ability to follow instructions under long text creation and professional constraints. To address challenges like incomplete descriptions and inconsistent terminology in real protocols, we extract key information from enhanced data and use Deepseek-R1 guo2025deepseek to automatically generate three tasks of different difficulty levels: Easy (atomic steps, no dependencies), Standard (two-level nesting, basic dependencies) and Difficult (three or more levels, complex dependencies). The examples generated by LLM must be strictly extracted from the original text rather than constructed. Up to five easy, three standard, and one difficult task are generated per protocol, retaining original step numbering. Priority is given to representative subtasks if source material is insufficient. Details on the format are in Appendix C.4.
@@
Protocol Reasoning (REA) task introduces Chain of Thought (CoT) prompting based on ERR and GEN tasks to probe explicit reasoning pathways regarding experimental intent and potential risks. For GEN task, we designed a structured CoT template comprising <Objective>, <Precondition>, <Phase>, <Prameter>and <Structure>. This data guides the model to sort out the experimental objectives and conditional dependencies, ensuring that the generated content is both scientifically reasonable and in line with the field format specifications. For ERR task, CoT instances describe identified error types and analyze potential experimental risks and safety hazards. Details on the CoT format are in Appendix C.5.
@@ 2.3 Multi-stage quality control
To guarantee data reliability and quality, we implemente a multi-stage quality control process with a three-phase automated self-filtering pipeline. First, automated checks validate structural integrity (e.g. correct step indices, answer–question alignment). Second, template-based and semantic-similarity tests remove any mismatches between generated and reference answers. Third, outputs for challenging cases are regenerated and the best versions selected. Following automated filtering, a randomly selected 10% of instances underwent manual review by an expert group comprising over 5 PhD-level biologists. This manual review focuse on critical aspects such as experimental safety constraints, accuracy of technical terminology, and the scientific validity of procedural steps and reasoning chains. This multi-stage approach effectively balances dataset scale with the high fidelity required for evaluating complex biological protocol comprehension and generation.
@@ 3 Evaluation Metrics
We evaluate model performance using a combination of standard and domain-specific metrics. Table 1 provides an overview of all metrics used across different tasks. Details on standard metrics and LLM-as-a-Judge consistency are provided in the Appendix F. Below, we detail the novel domain-specific metrics we introduce: keyword-based content metrics and embedding-based structural metrics.
@@ 4 Experiments
We conducted experiments with a diverse range of LLMs to provide a broad perspective on current capabilities for biological protocol tasks. To this end, we selected 12 state-of-the-art models, encompassing three main categories: Closed-Source LLMs (o3-mini openai2025o3mini , GPT-4o achiam2023gpt , GPT-4-turbo openai2023gpt4turbo , Claude-3.7-sonnet anthropic2025claude37sonnet , Gemini-2.0-flash google2024gemini2 , Gemini-2.5-pro-exp google2025gemini2_5_pro_exp ), Open-Source LLMs (Deepseek-R1 guo2025deepseek , DeepSeek-V3 liu2024deepseek , QwQ-32B qwen2025qwq32b , Qwen-2.5-72b-instruct qwen2024qwen2_5_72b ), and Domain-Specific Bio-LLMs (BioMedGPT-10B luo2023biomedgpt , BioMistral-7B labrak2024biomistral , BioMistral-7B-DARE labrak2024biomistral ). For evaluation, we constructed a held-out test set consisting of approximately 1,000 representative samples per task. Full implementation details are provided in Appendix E.
4.1 Performance on Protocol Question Answering (PQA)
@@
On PQA task, evaluated LLMs show varied proficiency in comprehending and extracting information from biological protocols (Figure 4a and Table 2). The top performance is achieved by the closed-source model Gemini-2.5-pro-exp (70.27%Acc.). However, leading open-source models, notably Deepseek-R1 (67.83%Acc.) and Deepseek-V3 (66.58%Acc.), demonstrate strong capabilities, surpassing several other closed-source models. This highlights the increasing viability of advanced open-source models for domain-specific comprehension tasks, significantly narrowing the performance gap with closed-source alternatives in this domain. In contrast, the dedicated domain-specific models, BioMedGPT-10B and BioMistral-7B, performed exceptionally poorly (20-21%Acc.), exhibiting very high failure rates (33-53%Failed). This suggests that their limited parameter count and primary optimization for general biomedical text are insufficient for the broader procedural and technical understanding required by biological protocols, a capability currently better captured by larger, generalist LLMs. Accuracy generally correlated positively with confidence calibration, as measured by the Brier Score (BS). The top performer, Gemini-2.5-pro-exp, showed the best calibration (0.1949 BS), closely followed by Deepseek-R1 (0.1965 BS). Conversely, models with lower accuracy, such as GPT-4-turbo (57.92% Acc.), exhibited poorer calibration (0.3403 BS). This consistency suggests that models with stronger comprehension capabilities also provide more reliable confidence estimates for their predictions.
4.2 Performance on Step Ordering (ORD)